# WKB structure in a scalar model of flat bands

### v1

@@ 1. Introduction
and the bands are the eigenvalues of $e^{-i\langle z,k\rangle}H(\alpha)e^{i\langle z,k\rangle}:H^{1}(\mathbb{C}/\Lambda;\mathbb{C}^{4})\to L^{2}(\mathbb{C}/\Lambda,\mathbb{C}^{4})$, $z,k\in\mathbb{C}$, considered as functions of $k$:
$\cdots\leq-E_{j}(\alpha,k)\leq\cdots\leq-E_{1}(\alpha,k)\leq 0\leq E_{1}(\alpha,k)\leq\cdots\leq E_{j}(\alpha,k)\leq\cdots$ (1.5)
(For the general theory of Bloch–Floquet band spectrum see [TaZw23, §5.3] and for an introduction in this context [Zw24, §3.1,3.2].) A flat band occurs when $E_{1}(\alpha,k)\equiv 0$ and that is equivalent to (1.3).
Local solutions to $D_{\rm{S}}(\alpha)u=0$ (or equivalently solution to $P(\alpha,0)u_{0}=0$) are given by local holomorphic sections of a rank-2 holomorphic (topologically trivial) vector bundle over $\mathbb{C}/\Lambda$. We denote it by $\mathscr{E}(\alpha)$ and it is defined in §3.6. The first theorem describes the connection between the structure of this vector bundle and the behaviour of bands:
@@
We refer to $m$ in the third statement of the theorem as the multiplicity of a magic $\alpha\in\mathcal{A}$. Although the model we consider does not seem to be physically relevant, Li–Yang [LiYa] proved that the same trichotomy appears in the case of twisted multiple layers of graphene – see [Ya23] for a mathematical presentation and references to the physics literature. The fact that double Dirac points can occur at a discrete set of angles was not present in the physics literature. The computations in [LiYa] suggest that Dirac points should occur when the angle of twisting between two bilayer graphene wafers if close to 0.44.
@@ 1.1. Quantisation condition for real magic $\alpha$’s
$\alpha_{j+1}-\alpha_{j}=\gamma+o(1),\ \ j\to+\infty,\ \ \gamma\simeq\tfrac{3}{2}.$ (1.6)
The more accurate computations made in [Be*22] suggest that $\gamma\simeq 1.515$. In the scalar model (2.4) with $V(z)=U_{\rm{BM}}(z)U_{\rm{BM}}(-z)$ where $U_{\rm{BM}}$ is given by (2.2) we numerically observe the following rule for real elements of $\mathcal{A}$: real magic $\alpha$’s are double (in the sense of Theorem 1), $\alpha_{2j-1}=\alpha_{2j}$, $j=1,2,\cdots$,
@@ 1.2. Nonvanishing potentials
In view of these complications we consider a yet simpler model in which one can use separation of variables and the complex WKB. For that we take $\Lambda:=\mathbb{Z}^{2}$ and $W(x,y)=W(x)$, $(x,y)\in\mathbb{R}^{2}$. Taking the Fourier modes in $y$ we obtain a family of operators depending $\zeta=2\pi mi+k$, where $m\in\mathbb{Z}$ is a Fourier mode in $y$ and $k$ is in the fundamental domain of the dual lattice $2\pi\mathbb{Z}^{2}$:
@@
Suppose that $W(x)$ in an entire function satisfying $W(x+1)=W(x)$. If there exists a closed smooth curve ${\gamma}:[0,1]\to\mathbb{C}/\mathbb{Z}$, $\gamma^{\prime}(t)\neq 0$, such that
$\operatorname{Im}{\gamma}^{\prime}(t)W({\gamma}(t))=0,\ \ W(\gamma(t))\neq 0,\ \ t\in[0,1],$ (1.14)
@@
$W(z)\neq 0,\ \ |\operatorname{Im}(e^{-i\arg W_{0}}W(z))|\leq\operatorname{Re}(e^{-i\arg W_{0}}W(z))\ \ \ \text{ for $|\operatorname{Im}z|\leq 1$.}$ (1.15)
For instance, if $f(z)$ is entire and $f(z+1)=f(z)$ then $W(z)=M+f(z)$ will satisfy (1.15) for $M\geq\sqrt{2}\max_{|\operatorname{Im}z|\leq 1}|f(z)|$.
@@
Acknowledgements. We would like to thank Zhongkai Tao for help with the material in §3.6, John Lott for the reference [Da07] and Mengxuan Yang for comments on the first version of this paper. The authors were partially supported by the Simons Targeted Grant Award No. 896630 “Moiré Materials Magic”, while the first named author was partially supported by NSF CAREER grant DMS-1749858.
@@ 2.1. Chiral model
Here we review some of its properties. It is derived from the the chiral model of twisted bilayer graphene (TBG) [TKV19], [Be*22], [BHZ24]:
@@
It is shown, together with the magic $\alpha$’s for the chiral model (2.1), both with $U$ given by (2.3), in Figure 4
@@ 3.1. Family of protected states
In view of the mapping properties in (2.24), we conclude that $\Pi(\alpha)=\mathbb{C}u(\alpha)\oplus\mathbb{C}v(\alpha)$, where
@@
Remark. For $\alpha\in\mathbb{R}\setminus\{0\}$, having a kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{-1}}$ is a natural possibility for another reason than the one presented before (3.10). It implies that $\ker H_{0}(\alpha)\neq 0$. Since $\ker H_{\ell}(0)=0$, $\ell\neq\pm 1$, the chiral symmetry shows that $\dim\ker H_{\ell}(\alpha)\equiv 0\mod 2$ and if $\dim\ker H(\alpha)>2$, then there exists a unique (in view of (3.9)) $\ell\neq\pm 1$ such that $\dim\ker H_{\ell}(\alpha)=2$ (For $\alpha\in\mathbb{R}\setminus\mathcal{A}$, we have $\dim\ker H(\alpha)\leq 4$, as otherwise $\mathscr{R}_{4}H(\alpha)=H(\alpha)\mathscr{R}_{4}$ would show that $\dim\ker D_{\rm{S}}(\alpha)>2$ contradicting (2.15).) Since $\mathscr{R}_{4}:\ker H_{\ell}(\alpha)\to H_{-\ell}(\alpha)$, that means that $\ell\equiv-\ell\mod 6$, that is $\ell=0$ or $\ell=3$. The case $\ell=0$ corresponds to a nontrivial kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{-1}}$ (see (2.24)). The case $\ell=3$ is excluded by the assumption that $\alpha\notin\mathcal{A}$. In fact, we would then have a non-trivial kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{2}}$. But that would mean that we would have $v$, with $D_{\rm{S}}(\alpha)v=0$, $\bar{\tau}v_{1}(\tau z),v_{2}(\tau z)]^{t}=[\tau^{2}v_{1}(z),\tau^{2}v_{2}(z)]^{t}$, that is $v_{1}(\tau z)=-v_{1}(z)$ and $v_{2}(\tau z)=\tau^{2}v_{2}(z)$. In particular $v(z)=0$ and the theta function argument (see [TKV19] and [Zw24, §6]) shows that $\alpha\in\mathcal{A}$ (see Proposition 2).
@@ 3.3. Zeros of $z\mapsto u(\alpha,z)$
$G_{0}(\zeta,\bar{\zeta}):=v_{1}(z,\bar{z})|_{z=z_{1}+\zeta},\ \ \ G_{j}(\zeta,\bar{\zeta}):=e^{-\frac{i}{2}(k\bar{z}+\bar{k}z)}u_{1}(z,\bar{z})|_{z=z_{1}+\zeta}.$
@@ 3.5. Local structure of solutions
The image of $F$, $X:=\{(z,\bar{z}):z\in\mathbb{C}\}$, is a totally real linear subspace (over $\mathbb{R}$) of $\mathbb{C}^{2}$ in the sense that $X\cap iX=\{0\}$. In particular, (fwe use the notation ${\rm neigh}_{\Omega}(\Omega_{0})$ for an open neighbourhood of $\Omega_{0}$ in $\Omega$, $\mathscr{O}(\Omega)$ denotes holomorphic functions in $\Omega$, $\zeta=(\zeta_{1},\zeta_{2}\in\mathbb{C}^{2}$)
@@
For a fixed $z$ we have the standard theory showing existence of unique solutions which are holomorphic in $w$. (In fact, we have a global solution in $w$ – see [Sj19, §7.2].) We now see that $\partial_{\bar{z}}\mathbf{u}_{j}$ solves the equation (3.27) but with the initial condition $\partial_{\bar{z}}\mathbf{u}_{j}(z,\bar{z}_{0})=0$. But that means that $w\mapsto\partial_{\bar{z}}\mathbf{u}_{j}(w,z)=0$. ∎
@@ 3.6. Rank-2 holomorphic vector bundle
Figure 6. On the left: contour plot of $z\mapsto\alpha^{-1}\log|u(\alpha,z)|$, $\alpha=19.9$, where to $(2D_{\bar{z}})^{2}-\alpha^{2}U(z)U(-z))u(\alpha,z)=0$ and $U$ is the Bistritzer–MacDonald potential (2.3). We see the structure of zeros and the regions of exponential decay predicted in [HiZw25]. On the right: the plot of $z\mapsto\alpha^{-1}\log|u(\alpha,z)|$ over the the interval connected the center to the edge of the hexagon, continued through the edge and then to the translated center (top) and over the interval connecting the centers and perpendicular to the edge of the hexagon. For an animated version with varying $\alpha$ see https://math.berkeley.edu/~zworski/scalar_protected.mp4.
@@ 4.1. Mechanism for the creation of zeros for the Bistritzer–MacDonald potential
For the scalar model numerical calculation show that the real elements of $\mathcal{A}$ have multiplicity 2. In view of the arguments presented in §3.4 this means that the protected state $v(\alpha)$, $D_{\rm{S}}(\alpha)v(\alpha)=0$ has simple zeros (in the sense of (3.20)) at the corners of the hexagon. Using (1.2) we see that
$v(\alpha,z)=\begin{pmatrix}\alpha^{-1}2D_{\bar{z}}u(\alpha,z)\\ u(\alpha,z)\end{pmatrix},\ \ \ ((2D_{\bar{z}})^{2}-\alpha^{2}V(z))u(\alpha,z)=0.$
@@
$v(w,\bar{w})=f_{0}(w,\alpha)A_{0}(\bar{w}-\alpha^{-\frac{2}{3}}cw^{2})+f_{2}(w,\alpha)A_{2}(\bar{w}-\alpha^{-\frac{2}{3}}cw^{2}),$ (4.8)
@@
$f_{0}(w,\alpha)=g_{0}(w^{3},\alpha),\ \ \ f_{2}(w,\alpha)=wg_{1}(w^{3},\alpha),\ \ \ g_{j}(\bar{\zeta},\alpha)=\overline{g_{j}(\zeta,\alpha)}.$ (4.9)
@@ 4.2. An example with a quantisation condition for magic alphas
We now considerUU given by (4.1). The properties (2.2) are inherited from the same properties being satisfied by $U_{\rm{BM}}$. Hence, the set of magic alphas, $\mathcal{A}$ is well defined both for the corresponding chiral and scalar models. Here we will consider the latter. What changes is the structure of the characteristic variety $q^{-1}(0)$, $q:=(2\bar{\zeta})^{2}-V(z)$. It consists of two Lagrangian tori with three high tangency common points – it is shown in Figure 10.
@@ 4.2.1. Approximate solution near the edge
The ansatz (4.20) is in remarkable with the protected state away from the $\pm\frac{1}{2}$ that is away from the center of the hexagon – see Figure 12. That is particularly striking on the edge continued (in a different fundamental domain) to the center – see Figure 11. We stress that a much finer analysis would be needed to justify this agreement. That has to involve matching the ansatz (4.20) with the rougher ansatz near the center (4.25).
@@ 5. A toy model
A much simpler version of (1.1) is given when $V(z)=W(z)^{2}$ and $W\neq 0$. In that case the characteristic variety $q^{-1}(0)$, $q=(2\bar{\zeta})^{2}-W(z)^{2}$ is given by a union of two disjoint tori $\{(x,\xi):2\bar{\zeta}=\pm W(z)\}$, $\zeta=\frac{1}{2}(\xi_{1}-i\xi_{2})$. These tori are Lagrangian if and only of $\operatorname{Im}\partial_{z}W(z)\equiv 0$ but we are interested in the general case of complex trigonometric polynomials $W(z)$ satisfying the condition
Here $\Lambda$ is a lattice of periodicity of $W$ (not necessarily the lattice we considered so far).
Figure 14. The characteristic variety $q^{-1}(0)$ for the classical symbol of (1.1), $q(x,\xi)=(2\bar{\zeta})^{2}-V(z)$ with $V(z)=W(z)^{2}$ and $W(z)\neq 0$. It is a union of two disjoint tori projecting diffeormorphically to the base torus. The tori are Lagrangian only if $\operatorname{Im}\partial_{z}W(z)\equiv 0$. Flat bands do not occur for this model but we can ask for the values of $\alpha$ for which $\ker_{H^{2}(\mathbb{C}/\Lambda)}P(\alpha,k)\neq\{0\}$. An answer in a special case is given in Theorem 3.

### v2

@@ 1. Introduction
and the bands are the eigenvalues of $e^{-i\langle z,k\rangle}H(\alpha)e^{i\langle z,k\rangle}:H^{1}(\mathbb{C}/\Lambda;\mathbb{C}^{4})\to L^{2}(\mathbb{C}/\Lambda,\mathbb{C}^{4})$, $z,k\in\mathbb{C}$, considered as functions of $k$ (they are $\pm$ the singular values of $D_{\rm{S}}(\alpha)+k$):
$\cdots\leq-E_{j}(\alpha,k)\leq\cdots\leq-E_{1}(\alpha,k)\leq E_{1}(\alpha,k)\leq\cdots\leq E_{j}(\alpha,k)\leq\cdots$ (1.5)
(For the general theory of Bloch–Floquet band spectrum see [TaZw23, §5.3] and for an introduction in this context [Zw24, §3.1,3.2].) A flat band occurs at $\alpha$ such that $E_{1}(\alpha,k)\equiv 0$ and that is equivalent to (1.3).
Local solutions to $D_{\rm{S}}(\alpha)u=0$ (or equivalently solutions to $P(\alpha,0)u_{0}=0$) are given by local holomorphic sections of a rank-2 holomorphic (topologically trivial) vector bundle over $\mathbb{C}/\Lambda$. We denote it by $\mathscr{E}(\alpha)$ and it is defined in §3.6. The first theorem describes the connection between the structure of this vector bundle and the behaviour of bands:
@@
We refer to $m$ in the third statement of the theorem as the multiplicity of a magic $\alpha\in\mathcal{A}$. Although the model we consider does not seem to be physically relevant, Li–Yang [LiYa] proved that the same trichotomy appears in the case of twisted multiple layers of graphene – see [Ya23] for a mathematical presentation and references to the physics literature. The fact that double Dirac points can occur at a discrete set of angles was not present in the physics literature. The computations in [LiYa] suggest that Dirac points should occur when the angle of twisting between two bilayer graphene wafers is close to 0.44.
@@ 1.1. Quantisation condition for real magic $\alpha$’s
$\alpha_{j+1}-\alpha_{j}=\gamma+o(1),\ \ j\to+\infty,\ \ \gamma\approx\tfrac{3}{2}.$ (1.6)
The more accurate computations made in [Be*22] suggest that $\gamma\approx 1.515$. In the scalar model (2.4) with $V(z)=U_{\rm{BM}}(z)U_{\rm{BM}}(-z)$ where $U_{\rm{BM}}$ is given by (2.2) we numerically observe the following rule for real elements of $\mathcal{A}$: real magic $\alpha$’s are double (in the sense of Theorem 1), $\alpha_{2j-1}=\alpha_{2j}$, $j=1,2,\cdots$,
@@ 1.2. Nonvanishing potentials
In view of these complications we consider a yet simpler model in which one can use separation of variables and complex WKB. For that we take $\Lambda:=\mathbb{Z}^{2}$ and $W(x,y)=W(x)$, $(x,y)\in\mathbb{R}^{2}$. Taking the Fourier modes in $y$ we obtain a family of operators depending on $\zeta=2\pi mi+k$, where $m\in\mathbb{Z}$ is a Fourier mode in $y$ and $k$ is in the fundamental domain of the dual lattice $2\pi\mathbb{Z}^{2}$:
@@
Suppose that $W(x)$ in an entire function satisfying $W(x+1)=W(x)$. Put $W_{0}:=\int_{0}^{1}W(x)\,dx$ and assume that $W_{0}\neq 0$. If there exists a closed smooth non-self-intersecting curve ${\gamma}:[0,1]\to\mathbb{C}/\mathbb{Z}$, $\gamma^{\prime}(t)\neq 0$, such that
$\operatorname{Im}{\gamma}^{\prime}(t)W_{0}^{-1}W({\gamma}(t))=0,\ \ W(\gamma(t))\neq 0,\ \ t\in[0,1],$ (1.14)
@@
$W(z)\neq 0,\ \ |\operatorname{Im}(W_{0}^{-1}W(z))|\leq\operatorname{Re}(W_{0}^{-1}W(z))\ \ \ \text{ for $|\operatorname{Im}z|\leq 1$.}$ (1.15)
For instance, if $f(z)$ is entire and $f(z+1)=f(z)$ and $\int_{0}^{1}f(z)dz=0$, then $W(z)=M+f(z)$ will satisfy (1.15) for $M\geq\sqrt{2}\max_{|\operatorname{Im}z|\leq 1}|f(z)|$.
@@ 2.1. Chiral model
The scalar model is derived from the the chiral model of twisted bilayer graphene (TBG) [TKV19], [Be*22], [BHZ24]:
@@
It is shown, together with the magic $\alpha$’s for the chiral model (2.1), both with $U$ given by (2.3), in Figure 4.
@@ 3.1. Family of protected states
In view of the mapping properties in (2.24), we conclude that $\operatorname{ran}\Pi(\alpha)=\mathbb{C}u(\alpha)\oplus\mathbb{C}v(\alpha)$, where
@@
Remark. For $\alpha\in\mathbb{R}\setminus\{0\}$, having a kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{-1}}$ is a natural possibility for another reason than the one presented before (3.10). It implies that $\ker H_{0}(\alpha)\neq 0$. Since $\ker H_{\ell}(0)=0$, $\ell\neq\pm 1$, the chiral symmetry shows that $\dim\ker H_{\ell}(\alpha)\equiv 0\mod 2$ and if $\dim\ker H(\alpha)>2$, then there exists a unique (in view of (3.9)) $\ell\neq\pm 1$ such that $\dim\ker H_{\ell}(\alpha)=2$ (For $\alpha\in\mathbb{R}\setminus\mathcal{A}$, we have $\dim\ker H(\alpha)\leq 4$, as otherwise $\mathscr{R}_{4}H(\alpha)=H(\alpha)\mathscr{R}_{4}$ would show that $\dim\ker D_{\rm{S}}(\alpha)>2$ contradicting (2.15).) Since $\mathscr{R}_{4}:\ker H_{\ell}(\alpha)\to H_{-\ell}(\alpha)$, that means that $\ell\equiv-\ell\mod 6$, that is $\ell=0$ or $\ell=3$. The case $\ell=0$ corresponds to a nontrivial kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{-1}}$ (see (2.24)). The case $\ell=3$ is excluded by the assumption that $\alpha\notin\mathcal{A}$. In fact, we would then have a non-trivial kernel of $D_{\rm{S}}(\alpha)|_{L^{2}_{2}}$. But that would mean that we would have $v$ with $D_{\rm{S}}(\alpha)v=0$, $\bar{\tau}v_{1}(\tau z),v_{2}(\tau z)]^{t}=[\tau^{2}v_{1}(z),\tau^{2}v_{2}(z)]^{t}$, that is $v_{1}(\tau z)=-v_{1}(z)$ and $v_{2}(\tau z)=\tau^{2}v_{2}(z)$. In particular $v(z)=0$ and the theta function argument (see [TKV19] and [Zw24, §6]) shows that $\alpha\in\mathcal{A}$ (see Proposition 2).
@@ 3.3. Zeros of $z\mapsto u(\alpha,z)$
$G_{0}(\zeta,\bar{\zeta}):=v_{1}(z,\bar{z})|_{z=z_{1}+\zeta},\ \ \ G_{j}(\zeta,\bar{\zeta}):=e^{-\frac{i}{2}(k\bar{z}+\bar{k}z)}u_{j}(z,\bar{z})|_{z=z_{1}+\zeta}.$
@@ 3.5. Local structure of solutions
The image of $F$, $X:=\{(z,\bar{z}):z\in\mathbb{C}\}$, is a totally real linear subspace (over $\mathbb{R}$) of $\mathbb{C}^{2}$ in the sense that $X\cap iX=\{0\}$. In particular, (we use the notation ${\rm neigh}_{\Omega}(\Omega_{0})$ for an open neighbourhood of $\Omega_{0}$ in $\Omega$, $\mathscr{O}(\Omega)$ denotes holomorphic functions in $\Omega$, $\zeta=(\zeta_{1},\zeta_{2}\in\mathbb{C}^{2}$)
@@
For a fixed $z$, the standard theory shows existence of unique solutions which are holomorphic in $w$. (If $w\mapsto V(w,z)$ is entire, we have a global solution in $w$ – see [Sj19, §7.2].) We now see that $\partial_{\bar{z}}\mathbf{u}_{j}$ solves the equation (3.27) but with the initial condition $\partial_{\bar{z}}\mathbf{u}_{j}(z,\bar{z}_{0})=0$. But that means that $w\mapsto\partial_{\bar{z}}\mathbf{u}_{j}(w,z)=0$. ∎
@@ 3.6. Rank-2 holomorphic vector bundle
Figure 6. On the left: contour plot of $z\mapsto\alpha^{-1}\log|u(\alpha,z)|$, $\alpha=20$, where to $(2D_{\bar{z}})^{2}-\alpha^{2}U(z)U(-z))u(\alpha,z)=0$ and $U$ is the Bistritzer–MacDonald potential (2.3). We see the structure of zeros and the regions of exponential decay predicted in [HiZw25]. On the right: the plot of $z\mapsto\alpha^{-1}\log|u(\alpha,z)|$ over the the interval connected the center to the edge of the hexagon, continued through the edge and then to the translated center (top) and over the interval connecting the centers and perpendicular to the edge of the hexagon. For an animated version with varying $\alpha$ see https://math.berkeley.edu/~zworski/scalar_protected.mp4.
@@ 4.1. Mechanism for the creation of zeros for the Bistritzer–MacDonald potential
For the scalar model numerical calculation show that the real elements of $\mathcal{A}$ have multiplicity 2. In view of the arguments presented in §3.4 this means that the protected state $\mathbf{v}(\alpha)$, $D_{\rm{S}}(\alpha)\mathbf{v}(\alpha)=0$ has simple zeros (in the sense of (3.20)) at the corners of the hexagon. Using (1.2) we see that
$\mathbf{v}(\alpha,z)=\begin{pmatrix}\alpha^{-1}2D_{\bar{z}}u(\alpha,z)\\ u(\alpha,z)\end{pmatrix},\ \ \ ((2D_{\bar{z}})^{2}-\alpha^{2}V(z))u(\alpha,z)=0.$
@@
$v(w,\bar{w})=f_{0}(w,\alpha)Ai_{0}(\bar{w}-\alpha^{-\frac{2}{3}}cw^{2})+f_{2}(w,\alpha)Ai_{2}(\bar{w}-\alpha^{-\frac{2}{3}}cw^{2}),$ (4.8)
@@
$f_{0}(w,\alpha)=g_{0}(w^{3},\alpha),\ \ \ f_{2}(w,\alpha)=wg_{2}(w^{3},\alpha),\ \ \ g_{j}(\bar{\zeta},\alpha)=\overline{g_{j}(\zeta,\alpha)}.$ (4.9)
@@ 4.2. An example with a quantisation condition for magic alphas
We now consider $U$ given by (4.1). The properties (2.2) are inherited from the same properties being satisfied by $U_{\rm{BM}}$. Hence, the set of magic alphas $\mathcal{A}$ is well defined both for the corresponding chiral and scalar models. Here we will consider the latter. What changes is the structure of the characteristic variety $q^{-1}(0)$, $q:=(2\bar{\zeta})^{2}-V(z)$. It consists of two Lagrangian tori with three high tangency common points – it is shown in Figure 10.
@@ 4.2.1. Approximate solution near the edge
The ansatz (4.20) is in remarkable agreement with the protected state away from the $\pm\frac{1}{2}$ that is away from the center of the hexagon – see Figure 12. That is particularly striking on the edge continued (in a different fundamental domain) to the center – see Figure 11. We stress that a much finer analysis would be needed to justify this agreement. That has to involve matching the ansatz (4.20) with the rougher ansatz near the center (4.25).
@@ 5. A toy model
We now turn our attention to the toy model $V(x,y)=W(x)^{2}$ featured in Theorem 3. Here $W$ is a 1-periodic function and we take the lattice of periodicity $\Lambda=\mathbb{Z}^{2}$.
5.1. Existence of protected states
Before studying the toy model, we consider the question of existence of protected states for operators of the form (1.1) where $V$ is a general potential, periodic with respect to some lattice $\Lambda\subset\mathbb{C}$, which no longer needs to satisfy the symmetries (2.5).